OpenStudy (anonymous):

Please help me!!!! MEDAL and Fan Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18 12x + 4y = -12 15x = -30 8x + 4y = -4 14x = -10 6x - 8y = -36 -6y = -42 6x - y = -15 3y = 9

OpenStudy (anonymous):

@CuteGirl903k @gelow

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

@smartgirl14 @smartkid2000 @Hollypettit @ayomike1 @AriPotta @ganeshie8 @KlOwNlOvE

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes did you want something

OpenStudy (anonymous):

help on the question

OpenStudy (anonymous):

are these the things you chose from 12x + 4y = -12 15x = -30 8x + 4y = -4 14x = -10 6x - 8y = -36 -6y = -42 6x - y = -15 3y = 9

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

no clue sorry

OpenStudy (anonymous):

i think 6x - 8y = -36 or -6y = -42

OpenStudy (anonymous):

so what does that mean?

OpenStudy (anonymous):

thats all one answer choice